# largesymbols"0D largesymbols"0D

### v1

@@ 4.3 Real-data estimates
Figure 3: Real-Data Estimates Across 14 Datasets
Taobao: iPad mini
Taobao: iPad 4
@@
Anonymous: laptop
Anonymous: vacuum
Anonymous: washer
@@
Our second set of tests are estimations of the search model entirely from real data – that is, both real attributes and real consumer choices. We no longer have a ground truth for $\boldsymbol{\theta}$ and cannot directly measure estimation accuracy. Therefore, we assess whether the pretrained NNE’s estimates are sensible, by comparing them with the estimates from the GHK-based SMLE that uses the very large $R=1000$.
Figure 3 shows the estimates across the 14 datasets. Each plot corresponds to one dataset. The horizontal axis lists the individual parameters in $\boldsymbol{\theta}$. The blue bars are estimates by SMLE and the orange bars are estimates by pretrained NNE. A 95% confidence interval is shown on each bar. Overall, the estimates by pretrained NNE and by SMLE are closely matched.
We omit reporting computation times here, because they are very similar to those reported in Table 4 for Monte Carlo studies. Across all the datasets, the pretrained NNE uses only a fraction of the computation time required by SMLE.
@@ 5 Conclusion
The ease of computation offered by pretrained NNE may facilitate new applications. One notable area is to integrate structural models with real-time algorithms, e.g., dynamic pricing, ranking algorithms, recommender systems, bandit experiments (e.g., Agrawal et al. 2019, Jain et al. 2024). Because structural models provide estimates of the latent underlying consumer preferences, these estimates can be fed into real-time algorithms to improve their decisions, e.g., what price to set, what items to recommend. Counterfactual analyses can be used to evaluate and provide feedback to the algorithms. However, structural estimation is time-consuming, which has made such applications difficult. Pretrained NNE offers a way to estimate structural models in real time, allowing more seamless integration of structural models into these algorithms.
The main intention of this paper is to explore the feasibility of pretrained estimators, rather than develop a comprehensive estimator for search models. This leaves areas where the pretrained NNE can be further improved. We note two potential areas. The first area is to increase the richness of the data and the structural model. The search model considered in this paper is a standard yet simple one. Future work may allow for attributes that simultaneously affect utility and search costs, or allow for richer unobserved heterogeneity in utility and search costs. These extensions will require us to find additional reduced-form patterns to help identify the more complex structural model.
A second area for improvement is to address potential endogeneity. For example, if researchers have an instrumental variable (IV) for price, how do we build a pretrained estimator that leverages the instrument? At a high level, addressing this problem involves two tasks. The first task is to generate training datasets that include instruments. The second task is to pass relevant information regarding the instrument to the neural net. These tasks are not straightforward. However, incorporating IV in NNE will be a significant and valuable extension, because we frequently need to address endogeneity in structural estimation applications.

### v2

@@ 4.3 Real-data estimates
Figure 3: Real-Data Estimates Across 12 Datasets
eCommerce: laptop
eCommerce: vacuum
eCommerce: washer
@@
Our second set of tests estimates the search model entirely from real data – that is, both real attributes and real consumer choices. We no longer have a ground truth for $\boldsymbol{\theta}$ and cannot directly measure estimation accuracy. Therefore, we assess whether the pretrained NNE’s estimates are sensible, by comparing them with the estimates from the GHK-based SMLE that uses the very large $R=1000$.
Figure 3 shows the estimates across the 12 datasets. Each plot corresponds to one dataset. The horizontal axis lists the individual parameters in $\boldsymbol{\theta}$. The blue bars are estimates by SMLE and the orange bars are estimates by pretrained NNE. A 95% confidence interval is shown on each bar. Overall, the estimates by pretrained NNE and by SMLE are closely matched.
We omit reporting computation times in Figure 3, because they are quantitatively similar to those reported in Table 4 for Monte Carlo studies. Across all the datasets, the pretrained NNE uses only a fraction of the computation time required by SMLE.
@@ 5 Conclusion
The ease of computation offered by pretrained NNE may facilitate new applications. One notable area is to integrate structural models with real-time algorithms, such as dynamic pricing, ranking algorithms, recommender systems, bandit experiments (e.g., Agrawal et al. 2019, Jain et al. 2024). Because structural models provide estimates of the latent underlying consumer preferences, these estimates can be fed into real-time algorithms to improve their decisions, e.g., what price to set, what items to recommend. Counterfactual analyses can be used to evaluate and provide feedback to the algorithms. However, structural estimation is time-consuming, which has made such applications difficult. Pretrained NNE offers a way to estimate structural models in real time, allowing more seamless integration of structural models into these algorithms.
The intention of this paper is to explore the feasibility of pretrained estimators, rather than develop a comprehensive estimator for search models. This leaves areas where the pretrained NNE can be further improved. We note two potential areas. The first area is to increase the richness of the structural model. The search model considered in this paper is a standard yet simple one. Future work may allow for attributes that simultaneously affect utility and search costs, or allow for richer unobserved heterogeneity in utility and search costs. These extensions will require us to find additional reduced-form patterns to help identify the more complex structural model.
A second area for improvement is to address potential endogeneity. For example, if researchers have an instrumental variable (IV) for price, how do we build a pretrained estimator that can leverage the instrument? At a high level, addressing this problem involves two tasks. The first task is to generate training datasets that include instruments. The second task is to pass relevant information regarding the instrument to the neural net. These tasks are not straightforward. However, incorporating IV in NNE will be a significant and valuable extension, because we frequently need to address endogeneity in structural estimation applications.